OpenStudy (henryblah):

How do you differentiate f(2x-y^2) wrt to x and y? I would like a method or link to a webpage explaining how to do it. It's probably chain rule, but I don't know how to apply it here.

OpenStudy (anonymous):

single variable or multivariable? and is it f(x,y) = 2x-y^2?

OpenStudy (henryblah):

The whole problem is u(x,y)=(e^-x)f(2x-y^2). And I need to find partial differentiation du/dx and dv/dx. So I think it's multivariable but not sure.

OpenStudy (anonymous):

yes, that's multivariable.

OpenStudy (anonymous):

what is f(x) or f(x,y) defined as?

OpenStudy (henryblah):

umm, it's a function from R to R and is differentiable.

OpenStudy (henryblah):

I have to prove it satisfies a u(x,y) satisfies a differential equation if that helps.

OpenStudy (anonymous):

I haven't done anything with differential equations in multiple dimensions yet, sorry

OpenStudy (henryblah):

No worries

OpenStudy (henryblah):

Here's the question if anyone is interested http://i.imgur.com/7QTnYN1.jpg

OpenStudy (anonymous):

f(x)'=2 f(y)'=-2y

OpenStudy (henryblah):

But why is that? As in how did you work it out

OpenStudy (anonymous):

oh, that's the multidimensional calculus answer to the differentiation assuming f(x,y) = 2x-y^2

hartnn (hartnn):

do you know chain rule ?

OpenStudy (henryblah):

Yes, but I don't see how it would work here

OpenStudy (anonymous):

when we differentiate with respect to 'x', then 'y' will be zero. and when we differentiate with repect to 'y', then 'x' will be consider as zero

hartnn (hartnn):

\(\Large \dfrac{\partial u }{\partial x}\) means differentiating u w.r.t 'x' and treating 'y' as CONSTANT.

hartnn (hartnn):

\(\Large \dfrac{\partial u }{\partial y}\) means differentiating u w.r.t 'y' and treating 'x' as CONSTANT.

OpenStudy (henryblah):

Yes, but how you differentiate something like f(2x) then? if you don't know what f is.

OpenStudy (henryblah):

oh hmm chain rule right?

hartnn (hartnn):

derivative of f(2x) would be f'(2x) d/dx (2x) simply put, chain rule.

hartnn (hartnn):

yes :)

OpenStudy (henryblah):

Ok, i'll try it and see if it works.

hartnn (hartnn):

\(\Large [f(2x-y^4)]' = f'(2x-y^4) [2x-y^4]' \)

OpenStudy (henryblah):

Hmm, that's the step that I think i'm getting wrong. I get du/dx to be -e^-xf(2x-y^4) + 2e^-xf'(2x). But I think that f'(2x) should be f'(2x-y^4)

hartnn (hartnn):

thats correct. -e^-xf(2x-y^4) + 2e^-xf'(2x-y^4)

OpenStudy (henryblah):

Oh right, I think it works!

OpenStudy (henryblah):

Thanks for the help.

hartnn (hartnn):

welcome ^_^